OpenStudy (love_to_love_you):

The equation below represents a function. f(x)=2x^2-9 What is the average rate of change of the function over the interval 3 <= x <= 6

OpenStudy (solomonzelman):

you just need to find the slope [of the secant] from the point wehre x=3, to the point where x=6.

OpenStudy (solomonzelman):

So, first, can you find \(f(3)\) and \(f(6)\) for me?

OpenStudy (love_to_love_you):

9 and 63?

OpenStudy (solomonzelman):

yes correct. f(3)=9 f(6)=63 Now, the slope between these 2 points is: \({\rm m}=\dfrac{f(6)-f(3)}{6-3}=\dfrac{63-9}{6-3}=?\)

OpenStudy (love_to_love_you):

54/3=18

OpenStudy (solomonzelman):

So the average rate of change on the interval [3,6] is 18

OpenStudy (solomonzelman):

any questions?

OpenStudy (love_to_love_you):

Could you help me with another question?

OpenStudy (solomonzelman):

yes, I think I can

OpenStudy (love_to_love_you):

Which of the following functions represents a geometric sequence? Why? a. The function f(x) = x^3 represents a geometric sequence because each term is cubed to make it greater. b. The function f(x) = 4x represents a geometric sequence because each term is the next higher multiple of 4. c. The function f(x) = 4x- 4 represents a geometric sequence because each term is 4 more than the previous term. d. The function f(x) = 4^x represents a geometric sequence because each term is 4 times as great as the previous term.

OpenStudy (love_to_love_you):

@SolomonZelman